# Grounding AI Explanations in Experience: A Reflective Cognitive Architecture for Clinical Decision Support

### v1

@@ A.4 Dataset Details
A.5 Explanation of Traditional MLs
@@
Unlike LLMs, traditional machine learning algorithms cannot directly produce textual explanations. Therefore, for Lasso Regression and CatBoost, we input the prediction results, along with the feature coefficients or feature importance produced by each method, into a Qwen3-235B, which performs best among all LLMs in explanation scores. Then It’s prompted to generate explanations based on prediction results and coefficients or importance.
A.6 LLM-based Agents
A.6.1 LLM+Tools
@@
We pre-defined a logistic regression function and a decision tree function to conduct relevant analysis,. These two functions can accept sample data and output feature correlation coefficients (for the logistic regression function) or feature importance (for the decision tree function) based on their built-in logic. For the LLM+tools method, when the LLM determines that tool invocation is necessary, it first organizes the data into a pre-specified format and sends it to the tool; after receiving the results returned by the tool, the LLM further uses these results to achieve data interpretation.
A.6.2 LLM+Code
@@
The code interpreter of o3-mini is one of its core functional modules, supporting dynamic tool generation and code execution. As a cost-effective inference model released by OpenAI, o3-mini exhibits excellent performance in STEM (Science, Technology, Engineering, Mathematics) fields—including science, mathematics, and coding (OpenAI, 2023). To this end, we input data into the LLM + code agent, and explicitly instruct the LLM to analyze the target problem by writing code; this setup aims to investigate whether the code interpreter can enhance the LLM’s understanding of the data.
A.7 Rubric of Explanation Criteria
Considering the value of clinical application, We designed detailed scoring rubric on each criterion for doctors to reference. Cognitive Load, Logical Argumentation, Evidence-Based Medicine and Cognitive Biasing are presented sequentially in Section
A.7.1 Cognitive Load(CL)
@@ A.7.2 Logical Argumentation(LA)
A.7.3 Evidence-Based Medicine(EBM)
@@ A.7.4 Cognitive Biasing(CB)
A.8 An Explanation Case
@@
A.9 Main Experiment Results
Table 3-5 present the explanation scores of all methods across the three datasets. As can be observed from the tables, the RCA methods outperform all baselines and achieve the highest scores across every dataset. This result indicates that the rules derived from in-depth understanding and summarization of data can effectively enhance the quality of interpretation. Among the four metrics, "Cognitive Biasing" consistently yields the lowest scores. This phenomenon suggests that when generating explanations, LLMs tend to prioritize listing factors that support their own conclusions—a characteristic that also aligns with the intuitive behavioral patterns of humans.
@@ A.9 Main Experiment Results
LLM-based
Method
Qwen2.5-72B 0.2698 -0.0073 0.1154 0.7470 0.4244 0.3333 0.3101 -0.0870 0.0851
@@
Reasoning
LLMs
DeepSeek-R1 0.1429 0.0805 0.1290 0.7229 0.3615 0.3200 0.2124 -0.1367 0.0741
@@
LLM-based
Agent
Qwen2.5-72B+Tools 0.2857 0.0023 0.1176 0.7470 0.4244 0.3333 0.3161 -0.0797 0.0851
@@
CRT CL LA EBM CB
Lasso 7.60 8.33 7.31 7.21
Traditional
MLs
Catboost 7.39 8.13 7.31 7.08
Qwen2.5-72B 7.92 8.16 8.25 7.14
DeepSeek-V3 7.98 8.27 8.22 7.19
DeepSeek-V3.1 7.41 8.11 8.13 6.73
LLM-Based
Methods
GPT-4.1 7.79 8.25 8.30 7.31
DeepSeek-R1 7.82 7.97 8.27 7.25
o3-mini 6.84 7.22 6.24 6.19
o4-mini 7.56 8.06 8.25 6.81
Qwen3-30B 7.87 8.49 8.41 7.06
Qwen3-235B 8.22 8.81 8.71 7.49
Reasoning
LLMs
GPT-5 7.63 8.22 8.40 7.05
Qwen2.5-72B+Tools 7.87 8.35 8.29 7.16
DeepSeek-V3.1+Tools 7.87 8.41 8.35 6.51
GPT-4.1+Tools 7.89 8.30 8.32 7.35
DeepSeek-R1+Tools 7.87 8.00 8.17 7.14
o3-mini+Tools 6.22 7.30 6.20 6.16
o4-mini+Tools 7.29 7.87 8.05 5.73
Qwen3-30B+Tools 7.79 8.63 8.41 6.57
Qwen3-235B+Tools 8.13 8.86 8.86 7.11
LLM-Based
Agents
o3-mini+Code 6.93 7.60 6.68 6.17
RCA + Qwen2.5 8.24 8.89 8.47 7.61
RCA RCA + GPT-4.1 8.16 8.59 8.87 7.62
@@ Methods
Diabetes CL LA EBM CB
Lasso 7.83 8.39 8.33 6.24
Traditional
MLs
Catboost 7.71 8.27 8.19 6.12
Qwen2.5-72B 7.61 8.20 8.10 5.98
DeepSeek-V3 7.55 8.04 8.29 5.99
DeepSeek-V3.1 7.85 8.23 8.12 6.04
LLM-Based
Methods
GPT-4.1 7.84 8.14 8.33 5.85
DeepSeek-R1 7.73 8.12 7.99 5.88
o3-mini 7.67 8.12 8.16 5.89
o4-mini 7.82 8.13 8.08 6.13
Qwen3-30B 7.94 8.43 8.51 6.72
Qwen3-235B 7.84 8.41 8.31 6.30
Reasoning
LLMs
GPT-5 7.66 8.24 8.53 6.24
Qwen2.5-72B+Tools 7.86 8.29 8.66 6.43
DeepSeek-V3+Tools 7.57 8.09 8.22 6.03
DeepSeek-V3.1+Tools 7.83 8.23 8.55 6.41
GPT-4.1+Tools 7.81 8.25 8.01 6.12
DeepSeek-R1+Tools 7.73 8.24 8.11 5.93
o3-mini+Tools 7.62 8.25 8.13 5.90
o4-mini+Tools 7.85 8.27 8.12 6.03
Qwen3-30B+Tools 7.82 8.13 8.08 6.13
Qwen3-235B+Tools 7.88 8.33 8.11 6.38
LLM-Based
Agents
o3-mini+Code 7.17 7.60 6.76 5.49
RCA + Qwen2.5 8.13 8.57 8.74 6.43
RCA RCA + GPT-4.1 8.03 8.38 8.63 6.94
@@
Heart Disease CL LA EBM CB
Lasso 7.60 8.27 8.63 5.94
Traditional
MLs
Catboost 7.50 8.21 8.72 5.94
Qwen2.5-72B 7.63 8.33 8.77 6.14
DeepSeek-V3 7.52 8.25 8.62 6.05
DeepSeek-V3.1 7.35 8.04 8.55 5.69
LLM-Based
Methods
GPT-4.1 7.45 8.08 8.64 5.77
DeepSeek-R1 7.40 7.98 8.59 5.75
o3-mini 7.33 7.89 8.84 5.46
o4-mini 7.31 7.83 8.73 5.38
Qwen3-30B 7.41 8.06 8.49 5.82
Qwen3-235B 7.67 8.34 8.84 6.36
Reasoning
LLMs
GPT-5 7.36 8.07 8.59 5.62
Qwen2.5-72B+Tools 7.63 8.34 8.74 6.0
DeepSeek-V3+Tools 7.43 8.21 8.55 6.11
DeepSeek-V3.1+Tools 7.37 7.97 8.46 5.72
GPT-4.1+Tools 7.42 8.01 8.58 5.70
DeepSeek-R1+Tools 7.38 7.87 8.55 5.73
o3-mini+Tools 7.31 7.90 8.76 5.48
o4-mini+Tools 7.35 7.96 8.51 5.66
Qwen3-30B+Tools 7.39 7.94 8.56 5.73
Qwen3-235B+Tools 7.58 8.28 8.91 6.20
LLM-Based
Agents
o3-mini+Code 7.28 7.88 8.67 5.38
RCA + Qwen2.5 7.62 8.47 8.94 6.18
RCA RCA + GPT-4.1 7.74 8.53 8.79 6.42
A.10 Robust Experiment Results
Table 6 presents the predictive performance across all datasets, while Tables 7-9 respectively show the explanation scores for each of the three datasets. As can be seen from the table data, RCA+GPT-4.1 achieves the best performance in nearly all predictive metrics while maintaining competitive explanation scores, with RCA+Qwen2.5 performing slightly worse in prediction task. Such minor fluctuations indicate that RCAenables LLMs to maintain robustness against data noise. Furthermore, a comprehensive analysis of the aforementioned tables reveals no direct correlation between predictive performance and explanation scores: for instance, Qwen3-235B, which performs poorly in terms of accuracy, achieves impressive scores across all four explanation metrics. This phenomenon demonstrates that modern LLMs possess strong capabilities in generating explanatory texts, regardless of whether their predictive results are correct or not.

### v2

@@ A.2 Dataset Details
A.3 Baselines
To comprehensively evaluate the performance of RCAin prediction and explanatory capability, we select a diverse range of baselines categorized into six types as follows. Details for Traditional MLs, Neural-Symbolic Networks and LLM-based Agents are provided in  A.3.1 and  A.3.2.
  • Traditional ML Models: We include Lasso regression (Tibshirani, 1996) and Catboost (Prokhorenkova et al., 2018). These models are standard for tabular data but produce statistical artifacts (e.g., coefficients, feature importance) rather than narrative explanations, representing a baseline for expert-driven interpretation. A ‘Qwen3-235B‘ is used to generate an explanation for these models.
  • Neural-symbolic Networks: To address a key alternative to our approach, we include two prominent neural-symbolic methods: Logic Tensor Networks (LTN)(Badreddine et al., 2022) and Logic Neural Networks (LNN) (Riegel et al., 2020). These baselines test whether explicit logic integration outperforms RCA’s emergent, LLM-based reasoning. We also used a ‘Qwen3-235B‘ to generate an explanation for these models
  • Medical LLM: We also include ‘Baichuan-M2’(Dou et al., 2025), a prominent large language model specifically optimized for the medical domain, to assess the performance of domain-specific models.
  • LLM-based Agents: We include two agent paradigms that reflect the state-of-the-art in LLM-driven analysis, comprising 9 and 1 methods, respectively. The LLM+Tools approach equips models with predefined functions, while the LLM+Code approach utilizes a code interpreter (OpenAI, 2023). These baselines test the hypothesis that tool use abstracts away the fine-grained data interaction necessary for deep understanding.
A.3.1 Explanation of Traditional MLs and Neural-symbolic Networks
@@
Unlike LLMs, traditional machine learning and neural-symbolic algorithms cannot directly produce textual explanations. Therefore, for Lasso Regression and CatBoost we input the prediction results along with the feature coefficients or feature importance produced by each method, into a Qwen3-235B, which performs best among all LLMs in explanation scores. Then It’s prompted to generate explanations based on prediction results and coefficients or importance. For our neural-symbolic networks, we input the final prediction and the truth values of its constituent logical predicates into a Qwen3-235B to generate a transparent, rule-based explanation.
A.3.2 LLM-based Agents
@@
  • LLM+Tools We pre-defined a logistic regression function and a decision tree function to conduct relevant analysis,. These two functions can accept sample data and output feature correlation coefficients (for the logistic regression function) or feature importance (for the decision tree function) based on their built-in logic. For the LLM+tools method, when the LLM determines that tool invocation is necessary, it first organizes the data into a pre-specified format and sends it to the tool; after receiving the results returned by the tool, the LLM further uses these results to achieve data interpretation.
@@
  • LLM+Code The code interpreter of o3-mini is one of its core functional modules, supporting dynamic tool generation and code execution. As a cost-effective inference model released by OpenAI, o3-mini exhibits excellent performance in STEM (Science, Technology, Engineering, Mathematics) fields—including science, mathematics, and coding (OpenAI, 2023). To this end, we input data into the LLM + code agent, and explicitly instruct the LLM to analyze the target problem by writing code; this setup aims to investigate whether the code interpreter can enhance the LLM’s understanding of the data.
A.4 Explanation Quality Metrics
As noted in  4.1, we developed four criteria grounded in cognitive science and medical practice: Cognitive Load (CL), Logical Argumentation (LA), Evidence-based Medicine (EBM), and Cognitive Biasing (CB). These criteria have been recognized by 3 doctors as clinically valuable, as they align with real-world medical assessment needs for evaluating the clarity, rationality, and reliability of explanatory content. We invited 3 doctors to score 100 samples for each of the core methods: Traditional MLs (‘CatBoost’), LLM-based Methods (‘GPT-4.1’), Reasoning LLMs (‘Qwen3-235B’), LLM-based Agents (‘Qwen3-235B + tools’), and RCA (‘RCA+GPT-4.1’). For all other methods, we used ‘Qwen3-30B’ with a carefully crafted prompt that achieved 90% scoring agreement with doctors. Then doctors and LLM scored each explanation on a scale of 1 to 10 for each criterion based on a detailed rubric we designed. The detailed rubrics are listed below.
A.4.1 Cognitive Load(CL)
@@ A.4.2 Logical Argumentation(LA)
A.4.3 Evidence-Based Medicine(EBM)
@@ A.4.4 Cognitive Biasing(CB)
A.5 EDGE CASE ANALYSIS: HANDLING EXTREME ANOMALIES
While our primary design objective for RCA was to generate high-fidelity, evidence-based explanations for the general patient population, the architecture is inherently well-suited to identifying and handling statistical anomalies. A critical challenge in real-world medical data is distinguishing between robust clinical signals and extreme noise (e.g., data entry errors or sensor malfunctions).
To demonstrate the robustness of RCA in such scenarios, we analyze a specific real-world case from the CRT dataset that contains extreme outlier values.
A.5.1 Case Description
The patient record presents with the following clinical features, including two physiologically improbable values:
  • Platelet Count: 2300.0 (Population Mean $\approx$ 230.0)
  • Hemoglobin: 1350.0 (Population Mean $\approx$ 118.0)
  • Other Features: CVC catheterization, undergoing chemotherapy, D-dimer 0.89, Age 64.
  • Ground Truth: No catheter-related thrombosis.
Baseline methods, lacking a grounded understanding of the global data distribution, often misinterpreted these extreme values as high-risk indicators due to their magnitude, leading to incorrect predictions of thrombosis.
A.5.2 RCA Performance and Analysis
In contrast, RCA correctly predicted "no catheter-related thrombosis." The distribution-aware rules check mechanism (§3.3) successfully utilized the global statistical context ($\mathcal{D}_{train}$) to flag these values as deviations falling far outside the 99th percentile.
The explanation generated by RCA explicitly articulates this reasoning:
"The patient has a moderately elevated D-dimer (0.89), a very high platelet count (2300.0) and an extremely high hemoglobin (1350.0), both of which are clear outliers likely due to data entry errors or extreme physiological abnormalities; such values far exceed physiologic ranges. … Laboratory outlier values alone should not be used as the primary basis for CRT prediction; this case warrants immediate clinical verification."
A.5.3 Discussion on Anomaly Handling
This case highlights a key safety feature of our framework. RCA does not attempt to algorithmically distinguish between a "genuine rare edge case" and "data noise/error," as both manifest as significant deviations from the learned distribution. Instead, it adopts a clinically responsible strategy:
  1. 1.
    Detection: Identifying the anomaly via the distribution check.
  2. 2.
    Logging: Explicitly noting the discrepancy in the explanation.
  3. 3.
    Flagging: Reducing the predictive weight of the outlier features and advising human review rather than forcing an unsupported high-confidence prediction.
This approach ensures that RCA remains robust to noise while providing necessary alerts for expert intervention.
A.6 Implicit Sub-Population Analysis via Distribution-Aware Reasoning
While RCA  does not include an explicit module for pre-defined sub-group analysis, its architecture enables it to organically identify and apply tailored logic to distinct sub-populations. This emergent capability stems from its core process of generating and verifying rules against the complete statistical context of the data. By weighing evidence within the global distribution, the model implicitly learns context-dependent rules that behave differently for different data regimes.
A clear illustration of this is found in the final rule base generated for the CRT dataset. The model effectively stratified patients into three distinct physiological sub-populations based on their D-dimer levels, a key biomarker for thrombosis. For each group, it learned a unique diagnostic standard, demonstrating a nuanced understanding that goes beyond simple thresholding. The sub-populations and their corresponding logic are as follows:
  • High-Risk Sub-population (D-dimer $\geq$ 2.0): For patients in this stratum, the model identified that the high D-dimer level is a strong intrinsic signal for CRT. Consequently, it learned a rule (Rule 3) requiring only one additional moderate risk factor (e.g., chemotherapy) to confirm a positive diagnosis. The model correctly inferred that the baseline risk in this group is already high, lowering the evidence bar for a final prediction.
  • Intermediate-Risk Sub-population (D-dimer 1.0–1.99): In this group, the model recognized a greater degree of ambiguity, as the D-dimer level is elevated but not definitive. As a result, it learned a more conservative rule (Rule 4) that necessitates a higher burden of proof: at least two moderate risk factors must be present to predict CRT. This reflects an understanding that more corroborating evidence is needed to reach a conclusion in this ambiguous zone.
  • Low-Risk Sub-population (D-dimer << 1.0): For these patients, the model treated the low D-dimer level as strong contra-evidence for CRT. To overcome this negative signal, it learned its strictest rule (Rule 5), which requires the accumulation of three or more moderate risk factors before making a positive prediction. This demonstrates the model’s ability to demand a high volume of evidence to overturn a strong baseline indication.
As these rules demonstrate, RCA  does not treat D-dimer as a biomarker with a single, universal threshold. Instead, by grounding its reasoning in the complete data distribution, it successfully identified distinct risk profiles and learned context-dependent diagnostic standards for each sub-population. This ability to implicitly stratify a population and apply tailored logic is a key feature of its data-driven approach, enhancing both the accuracy and the interpretability of its conclusions.
A.7 An Explanation Case
@@
Neural-symbolic Networks LTN 0.5161 0.0085 0.1176 0.7635 0.5247 0.6946 0.6010 -0.0386 0.2222
LNN 0.2581 0.1287 0.1481 0.7437 0.4974 0.6787 0.5389 -0.0598 0.2393
LLM-based Method Qwen2.5-72B 0.2698 -0.0073 0.1154 0.7470 0.4244 0.3333 0.3101 -0.0870 0.0851
@@
Reasoning LLMs DeepSeek-R1 0.1429 0.0805 0.1290 0.7229 0.3615 0.3200 0.2124 -0.1367 0.0741
@@
Medical LLM Baichuan-M2 0.1587 0.0845 0.1311 0.6988 0.2772 0.4186 0.0725 -0.7779 0.1095
LLM-based Agent Qwen2.5-72B+Tools 0.2857 0.0023 0.1176 0.7470 0.4244 0.3333 0.3161 -0.0797 0.0851
@@
Metrics CL LA EBM CB
Algorithms Mean. Var. Mean. Var. Mean. Var. Mean. Var.
Traditional MLs Lasso 7.60 0.39 8.33 0.36 7.31 0.63 7.21 0.56
Catboost 7.39 0.50 8.13 0.53 7.31 0.78 7.08 0.71
Neural-symbolic Networks LTN 7.47 0.23 8.25 0.32 8.05 0.61 7.27 0.59
LNN 7.51 0.28 8.35 0.33 7.63 0.74 7.06 0.54
LLM-based Methods Qwen2.5-72B 7.92 0.96 8.16 1.23 8.25 1.31 7.14 1.07
DeepSeek-V3 7.98 0.68 8.27 0.73 8.22 0.83 7.19 0.70
DeepSeek-V3.1 7.41 0.89 8.11 1.02 8.13 1.28 6.73 1.02
GPT-4.1 7.79 0.69 8.25 0.72 8.30 1.13 7.31 0.88
Reasoning LLMs DeepSeek-R1 7.82 1.03 7.97 1.29 8.27 1.60 7.25 1.07
o3-mini 6.84 0.96 7.22 1.23 6.24 1.62 6.19 1.21
o4-mini 7.56 1.18 8.06 1.36 8.25 1.80 6.81 1.49
Qwen3-30B 7.87 0.97 8.49 1.06 8.41 1.28 7.06 1.19
Qwen3-235B 8.22 0.94 8.81 1.15 8.71 1.25 7.49 1.08
GPT-5 7.63 0.97 8.22 1.06 8.40 1.32 7.05 1.05
Medical LLM Baichuan-M2 7.82 0.54 8.38 0.52 8.32 0.75 6.96 0.55
LLM-Based Agents Qwen2.5-72B+Tools 7.87 0.75 8.35 0.92 8.29 1.18 7.16 0.95
DeepSeek-V3.1+Tools 7.87 0.69 8.41 0.93 8.35 1.22 6.51 1.01
GPT-4.1+Tools 7.89 0.48 8.30 0.54 8.32 0.76 7.35 0.67
DeepSeek-R1+Tools 7.87 1.03 8.00 1.29 8.17 1.60 7.14 1.07
o3-mini+Tools 6.22 0.91 7.30 1.07 6.20 1.31 6.16 1.15
o4-mini+Tools 7.29 1.05 7.87 1.21 8.05 1.60 5.73 0.59
Qwen3-30B+Tools 7.79 0.86 8.63 0.99 8.41 1.13 6.57 0.83
Qwen3-235B+Tools 8.13 0.88 8.86 0.84 8.86 1.22 7.11 0.94
o3-mini+Code 6.93 1.04 7.60 1.19 6.68 1.57 6.17 1.23
RCA RCA+Qwen2.5 8.24 0.42 8.89 0.54 8.47 0.92 7.61 0.74
RCA+GPT-4.1 8.16 0.34 8.59 0.34 8.87 0.76 7.62 0.56
@@
Metrics CL LA EBM CB
Algorithms Mean. Var. Mean. Var. Mean. Var. Mean. Var.
Traditional MLs Lasso 7.83 0.32 8.39 0.26 8.33 0.30 6.24 0.40
Catboost 7.71 0.30 8.27 0.24 8.19 0.28 6.12 0.36
Neural-symbolic Networks LTN 7.68 0.35 8.22 0.29 8.28 0.33 6.18 0.43
LNN 7.82 0.33 8.35 0.27 8.45 0.31 6.28 0.39
LLM-based Methods Qwen2.5-72B 7.61 0.92 8.20 0.80 8.10 0.87 5.98 1.02
DeepSeek-V3 7.55 0.99 8.04 0.87 8.29 0.94 5.99 1.09
DeepSeek-V3.1 7.85 0.95 8.23 0.83 8.12 0.90 6.04 1.05
GPT-4.1 7.84 0.95 8.14 0.83 8.33 0.90 5.85 1.05
Reasoning LLMs DeepSeek-R1 7.73 1.02 8.12 0.90 7.99 0.97 5.88 1.12
o3-mini 7.67 1.05 8.12 0.93 8.16 1.00 5.89 1.15
o4-mini 7.82 1.00 8.13 0.88 8.08 0.95 6.13 1.10
Qwen3-30B 7.94 0.98 8.43 0.86 8.51 0.93 6.72 1.08
Qwen3-235B 7.84 0.89 8.41 0.77 8.31 0.84 6.30 0.99
GPT-5 7.66 1.05 8.24 0.93 8.53 1.00 6.24 1.15
Medical LLM Baichuan-M2 7.79 0.40 8.18 0.33 8.52 0.37 6.05 0.46
LLM-Based Agents Qwen2.5-72B+Tools 7.86 0.55 8.29 0.43 8.66 0.50 6.43 0.65
DeepSeek-V3.1+Tools 7.83 0.62 8.23 0.50 8.55 0.56 6.41 0.72
GPT-4.1+Tools 7.81 0.56 8.25 0.44 8.01 0.51 6.12 0.66
DeepSeek-R1+Tools 7.73 0.75 8.24 0.63 8.11 0.70 5.93 0.82
o3-mini+Tools 7.62 0.79 8.25 0.66 8.13 0.73 5.90 0.86
o4-mini+Tools 7.85 0.82 8.27 0.69 8.12 0.76 6.03 0.89
Qwen3-30B+Tools 7.82 0.73 8.13 0.61 8.08 0.68 6.13 0.80
Qwen3-235B+Tools 7.88 0.69 8.33 0.57 8.11 0.64 6.38 0.76
o3-mini+Code 7.17 1.09 7.60 0.97 6.76 1.04 5.49 1.19
RCA RCA+Qwen2.5 8.13 0.36 8.57 0.51 8.74 0.54 6.43 0.43
RCA+GPT-4.1 8.03 0.43 8.38 0.49 8.63 0.62 6.94 0.40
@@
Metrics CL LA EBM CB
Algorithms Mean. Var. Mean. Var. Mean. Var. Mean. Var.
Traditional MLs Lasso 7.60 0.35 8.27 0.28 8.63 0.32 5.94 0.42
Catboost 7.50 0.33 8.21 0.25 8.72 0.29 5.94 0.38
Neural-symbolic Networks LTN 7.45 0.38 8.15 0.31 8.35 0.35 6.20 0.45
LNN 7.65 0.36 8.32 0.29 8.58 0.33 6.35 0.41
LLM-based Methods Qwen2.5-72B 7.63 0.95 8.33 0.82 8.77 0.89 6.14 1.05
DeepSeek-V3 7.52 1.02 8.25 0.89 8.62 0.96 6.05 1.12
DeepSeek-V3.1 7.35 1.02 8.04 0.89 8.55 0.96 5.69 1.12
GPT-4.1 7.45 0.98 8.08 0.85 8.64 0.92 5.77 1.08
Reasoning LLMs DeepSeek-R1 7.40 1.05 7.98 0.92 8.59 0.99 5.75 1.15
o3-mini 7.33 1.08 7.89 0.95 8.84 1.02 5.46 1.18
o4-mini 7.31 1.03 7.83 0.90 8.73 0.97 5.38 1.13
Qwen3-30B 7.41 1.01 8.06 0.88 8.49 0.95 5.82 1.11
Qwen3-235B 7.67 0.92 8.34 0.79 8.84 0.86 6.36 1.02
GPT-5 7.36 1.08 8.07 0.95 8.59 1.02 5.62 1.18
Medical LLM Baichuan-M2 7.28 0.42 8.08 0.35 8.25 0.39 5.85 0.48
LLM-Based Agents Qwen2.5-72B+Tools 7.63 0.58 8.34 0.45 8.74 0.52 6.00 0.68
DeepSeek-V3.1+Tools 7.37 0.65 7.97 0.52 8.46 0.58 5.72 0.75
GPT-4.1+Tools 7.42 0.59 8.01 0.46 8.58 0.53 5.70 0.69
DeepSeek-R1+Tools 7.38 0.78 7.87 0.65 8.55 0.72 5.73 0.85
o3-mini+Tools 7.31 0.82 7.90 0.68 8.76 0.75 5.48 0.89
o4-mini+Tools 7.35 0.85 7.96 0.71 8.51 0.78 5.66 0.92
Qwen3-30B+Tools 7.39 0.76 7.94 0.63 8.56 0.70 5.73 0.83
Qwen3-235B+Tools 7.58 0.72 8.28 0.59 8.91 0.66 6.20 0.79
o3-mini+Code 7.28 1.02 7.88 0.99 8.67 1.06 5.38 1.12
RCA RCA+Qwen2.5 7.62 0.28 8.47 0.33 8.94 0.49 6.18 0.68
RCA+GPT-4.1 7.74 0.35 8.53 0.42 8.79 0.55 6.42 0.53
Table 6: Expert Preference for RCA. The values represent the percentage of times clinicians (N=5) chose RCA’s explanation as superior to the baseline’s.
Preference for RCA CRT Diabetes Heart Disease
Cognitive Load(CL) 82.7% 81.3% 84.7%
Logical Argumentation(LA) 78.0% 75.3% 72.0%
Evidence-Based Medicine(EBM) 72.7% 74.7% 73.3%
Cognitive Biasing(CB) 86.7% 89.3% 86.0%
A.8 Main Experiment Results
Tables 35 present a detailed evaluation of explanation quality across all methods on the CRT, Diabetes, and Heart Disease datasets, respectively. The results offer two crucial insights into the performance of our RCA framework.
First, an analysis of the mean scores demonstrates that RCA-based methods consistently achieve the highest average quality. Across all three datasets, RCA+Qwen2.5 and RCA+GPT-4.1 outperform all baselines, securing top ranks in Cognitive Load (CL), Logical Argumentation (LA), Evidence-Based Medicine (EBM), and Cognitive Biasing (CB). This indicates that, on average, the explanations generated by RCA are judged by clinical experts to be clearer, more logical, and more reliable.
Second, the variance data reveals a critical advantage in consistency. The RCA  framework consistently exhibits significantly lower variance compared to other LLM-based approaches, particularly the standalone Reasoning LLMs. For instance, in Table 3, the variance for RCA’s CL and LA scores (e.g., 0.3-0.5) is often two to three times lower than that of the baseline LLMs (e.g., 0.9-1.3). This low variance signifies a high degree of inter-rater agreement and confirms that RCA’s explanation quality is not just high, but also stable and predictable. In contrast, the high variance of standalone LLMs suggests their output quality is erratic, making them less trustworthy for critical applications. This dual achievement of delivering both the highest mean quality and the greatest consistency underscores the effectiveness of RCA’s structured, data-grounded reasoning process.
To further validate these quantitative findings from a user-centric perspective, we conducted a head-to-head preference study, asking clinicians to directly choose the better explanation between RCA  and the baselines. As shown in Table 6, the results provide a resounding endorsement of our framework. Across all datasets and criteria, clinicians overwhelmingly preferred RCA’s explanations, with preference rates consistently exceeding 70% and often surpassing 80%. Notably, the highest preference for RCA  was observed in Cognitive Biasing (CB), with rates reaching up to 89.3%. This is particularly significant because it shows that while generating a perfectly balanced argument is difficult for all models, clinicians find RCA’s structured approach to be substantially more balanced and less biased compared to the unstructured outputs of other LLMs. This direct preference data serves as powerful, qualitative proof of the practical superiority of our framework.
Finally, a notable cross-model trend is that scores for Cognitive Biasing (CB) are generally the lowest among the four metrics for nearly all models. This suggests a common tendency for LLMs to generate one-sided justifications, highlighting the inherent difficulty of producing truly balanced clinical arguments.